# Little Red Dots as Globular Clusters in Formation

### v1

@@ I Introduction
The discovery of “Black Hole Stars” (Naidu et al., 2025; de Graaff et al., 2025b), which resemble LRDs without significant UV continua, further highlights their unusual properties. These objects suggest the characteristic LRD V-shape can be spectroscopically decomposed into two distinct components: one in the rest-frame UV and one in the rest-frame optical (Sun et al., 2026). The rest-frame UV is well-described by a young stellar population, while the rest-frame optical is well-fit by a modified blackbody with temperatures 2000-7000 K and luminosities 109-11 L⊙ (de Graaff et al., 2025a). The spectral inflection occurs due to the relative strength of the UV and optical components, shifting as red as Å (de Graaff et al., 2025a).
@@
Here, we explore an alternative scenario in which LRDs are globular clusters in formation (e.g., Gieles et al., 2018). In this scenario, the LRD V-shape arises from an ultra-young stellar population producing the rest-frame UV continuum, while a supermassive star (SMS) produces the rest-frame optical. Young stellar clusters have high stellar multiplicity and suffer frequent dynamical interactions (Offner et al., 2023) that can drive runaway stellar mergers to form SMSs, even in relatively metal-rich environments (Gieles et al., 2018; Vergara et al., 2025). Thus, SMS formation may naturally occur in dense stellar environments to resolve both the globular cluster mass-budget problem (Renzini et al., 2015) and the origin of multiple populations (Carretta et al., 2010; Piotto et al., 2015; Gratton et al., 2010; Bastian and Lardo, 2018; Milone and Marino, 2022).
@@ II.1 Potential Supermassive Star Properties
where is the gravitational constant, is the proton mass, is the speed of light, and is the Thomson scattering cross section. Electron scattering is unlikely to be the dominant opacity at the observed , but H- opacity can be similar to or larger than electron scattering depending on the temperature (Gray, 2005; Kippenhahn et al., 2013). LRDs with will launch electron-scattering-driven winds. The SMS relation is estimated by equating to the Stefan-Boltzmann law as

### v2

@@ I Introduction
The discovery of “Black Hole Stars” (Naidu et al., 2025; de Graaff et al., 2025b), which resemble LRDs without significant UV continua, further highlights their unusual properties. These objects suggest the characteristic LRD V-shape can be spectroscopically decomposed into two distinct components: one in the rest-frame UV and one in the rest-frame optical (Barro et al., 2026; Sun et al., 2026). The rest-frame UV is well-described by a young stellar population, while the rest-frame optical is well-fit by a modified blackbody with temperatures 2000-7000 K and luminosities 109-11 L⊙ (de Graaff et al., 2025a). The spectral inflection occurs due to the relative strength of the UV and optical components, shifting as red as Å (de Graaff et al., 2025a).
@@
Here, we explore an alternative scenario in which LRDs are globular clusters in formation (e.g., Gieles et al., 2018). In this scenario, the LRD V-shape arises from an ultra-young stellar population producing the rest-frame UV continuum, while a supermassive star (SMS) produces the rest-frame optical. Young stellar clusters are extremely dense (Adamo et al., 2024; Claeyssens et al., 2023, 2026) with high stellar multiplicity that suffer frequent dynamical interactions (Offner et al., 2023) that can drive runaway stellar mergers to form SMSs, even in relatively metal-rich environments (Gieles et al., 2018; Vergara et al., 2025). Thus, SMS formation may naturally occur in dense stellar environments to resolve both the globular cluster mass-budget problem (Renzini et al., 2015) and the origin of multiple populations (Carretta et al., 2010; Piotto et al., 2015; Gratton et al., 2010; Bastian and Lardo, 2018; Milone and Marino, 2022).
@@ II.1 Potential Supermassive Star Properties
where is the gravitational constant, is the proton mass, is the speed of light, and is the Thomson scattering cross section. Electron scattering is likely to be the dominant opacity interior to the photosphere, and H- opacity can dominate at the cooler temperatures exterior to the photosphere (Gray, 2005; Kippenhahn et al., 2013). LRDs with will launch electron-scattering-driven winds. The SMS relation is estimated by equating to the Stefan-Boltzmann law as